# Metric Graph Kernels via the Tropical Torelli Map

### v1

@@ Abstract
We propose new graph kernels grounded in the study of metric graphs via tropical algebraic geometry. In contrast to conventional graph kernels that are based on graph combinatorics such as nodes, edges, and subgraphs, our graph kernels are purely based on the geometry and topology of the underlying metric space. A key characterizing property of our construction is its invariance under edge subdivision, making the kernels intrinsically well-suited for comparing graphs that represent different underlying spaces. We develop efficient algorithms for computing these kernels and analyze their complexity, showing that it depends primarily on the genus of the input graphs. Empirically, our kernels outperform existing methods in label-free settings, as demonstrated on both synthetic and real-world benchmark datasets. We further highlight their practical utility through an urban road network classification task.
@@ 1 Introduction
Graph kernels are powerful tools for measuring similarities between graphs and networks and have become essential in many machine learning tasks. Traditional graph kernels typically rely on combinatorial aspects of graphs, such as nodes, edges, and subgraphs, with additional information from node and edge labels and attributes [15, 36, 4, 21, 19]. While such kernels have demonstrated success in many domains, they can be limited in scenarios where graphs arise from geometric data, and where labels or attributes are absent or incomparable.
@@
We study metric graphs, which are geometric realizations of graphs with a length function on their edges. Specifically, we leverage the known one-to-one correspondence between metric graphs and abstract tropical curves from tropical algebraic geometry [9, 7]. This correspondence has been studied in tropical geometry via the tropical Torelli map [8, 6], which sends any metric graph to a flat torus. In computational settings that are especially relevant for machine learning, for a weighted graph with generic length function, we can compute a unique symmetric positive definite (SPD) matrix to represent the flat torus. This effectively allows us to identify weighted graphs with SPD matrices, which contain the intrinsic geometric and topological information on the underlying metric graph, using the tropical Torelli map.
@@
The space of SPD matrices plays a central role in information geometry and can be endowed with various Riemannian metrics to capture its rich geometric structure [16, 38, 33]. Motivated by insights from both tropical geometry and information geometry, we leverage the Bures–Wasserstein distance on the space of postive semi-definite (PSD) matrices and propose the tropical Torelli–Wasserstein (TTW) kernel. For simplicity and efficient computation, we also propose the tropical Torelli–Euclidean (TTE) kernel based on the Euclidean distance. Our graph kernels are purely based on the geometry and topology of the underlying metric graphs, thus are well-suited for comparing graphs that represent different underlying spaces.
@@
Contributions. Our work is the first to merge tropical geometry and information geometry in the context of machine learning, which proves to be a powerful combination. Our specific contributions are listed as follows:
@@
  • We propose new graph kernels leveraging the tropical Torelli map from tropical geometry. Our graph kernels are invariant under edge subdivisions and naturally extend to well-defined kernels on the underlying space of metric graphs.
@@
  • We develop algorithms to compute our graph kernels. We show that the computational complexity is dominated by graph genus—an algebraic topological measure of graph cycles. We give concrete analyses on different levels of graph sparsity.
  • We systematically compare our graph kernels with other label-free graph kernels on benchmark datasets. We show that our graph kernel outperforms other graph kernels in the absence of node/edge labels.
  • We conduct experiments on synthetic and real-world datasets. On synthetic datasets, we test the computational runtime of our graph kernels and show consistency with our theoretical analysis. For real-world data, we extract local urban road networks and apply our graph kernels to a classification task. Our results demonstrate the superior performance of our method over other existing graph kernels.
@@ 2 From Graphs to Metric Graphs
A graph consists of discrete sets of nodes and edges. A metric graph is a 1-dimensional metric space that can be realized as the underlying space of a graph. Any graph-based quantity extends well to metric graphs if it is compatible with edge subdivisions—a concept we will now introduce.
Edge Subdivisions. Let $G=(V,E)$ be a graph and $\ell:E\to\mathbb{R}_{+}$ be a length function. An edge subdivision of $e=[u,v]\in E$ is an operation on $G$ defined as follows: First, add a new node $w$ to $V$, and then replace $e$ by two edges $e^{\prime}=[u,w]$ and $e^{\prime\prime}=[w,v]$ such that $\ell(e)=\ell(e^{\prime})+\ell(e^{\prime\prime})$. An edge subdivision increases both the number of nodes and number of edges by 1. We say $G^{\prime}$ is a refinement of $G$, denoted as $G^{\prime}\geq G$, if $G^{\prime}$ can be obtained from $G$ by a sequence of edge subdivisions. By definition, all refinements of $G$ share the same underlying metric graph $|G|$.
We remark here that refinement introduces a new relation between graphs that differs from the usual notion of subgraph inclusion. Specifically, if $G^{\prime}$ is a refinement of $G$, then there exists an injection $V(G)\to V(G^{\prime})$ and a surjection $E(G^{\prime})\to E(G)$. In contrast, if $G$ is a subgraph of $G^{\prime}$, then $V(G)\to V(G^{\prime})$ and $E(G)\to E(G^{\prime})$ are both inclusions.
Definition 1.
A graph kernel $k$ is called a metric graph kernel if $k(G_{1}^{\prime},G_{2}^{\prime})=k(G_{1},G_{2})$ for any refinements $G_{1}^{\prime}\geq G_{1}$ and $G_{2}^{\prime}\geq G_{2}$, that is, $k$ is invariant under graph refinements.
Existing graph kernels are often based on nodes, edges, and subgraphs, which are not invariant under refinements. Therefore, they fail to be metric graph kernels.
@@
3 The Tropical Torelli Map for Weighted Graphs
In this section, we introduce the tropical Torelli map for weighted graphs; this is the map we use to send a graph to a matrix. We also provide an algorithm for its practical computation, where the uniqueness of the resulting matrix representation is required but challenging to achieve. Under the generic assumption on length functions, our algorithm outputs a unique SPD matrix for any weighted graph. For further details, including related background in tropical geometry and the original definition of the tropical Torelli map on metric graphs, see Appendix A.1.
A concise mathematical formulation of the process of mapping a metric graph to an SPD matrix is as follows: Let $G=(V,E)$ be a graph with a positive weight function $\ell:E\to\mathbb{R}_{+}$ which in our context should be interpreted as the length function on edges. The geometric realization of $G$, denoted as $|G|$, is a 1-dimensional metric space constructed by identifying each edge $e$ as the interval $[0,\ell(e)]$, and gluing all such intervals at endpoints. The geometric realization $|G|$ is also known as a metric graph, which is in one-to-one correspondence with abstract tropical curves in tropical geometry [9, 7]. The classical tropical Torelli map sends any metric graph to a flat torus or its tropical Jacobian variety. For a weighted graph with generic length function, we compute a unique SPD matrix $Q$ to represent the flat torus. Therefore we adapt the tropical Torelli map to weighted graphs by sending any $G$ to the SPD matrix $Q(G)$. We show that $Q(G)=Q(G^{\prime})$ for any refinement $G^{\prime}$ of $G$. Hence $Q(G)$ contains the intrinsic geometric and topological information on the underlying metric graph $|G|$.
@@ 3.1 Graph Homology
A concept that is central to our construction is the 1-homology group of a graph. The concept arises from algebraic topology and intuitively corresponds to cycles in a graph. In the following, we give mathematical details on the construction of the 1-homology group of a graph and how we obtain a matrix representation of the graph in terms of 1-cycles via the tropical Torelli map.
We first outline the 1-homology group as follows. Let $G=(V,E)$ be a connected graph with $n$ nodes and $m$ edges. Let $C_{0}(G;\mathbb{R})$ be the $n$-dimensional vector space spanned by $V$ and $C_{1}(G;\mathbb{R})$ be the $m$-dimensional vector space spanned by $E$. Fix an arbitrary orientation on $G$. The boundary map $\partial:C_{1}(G;\mathbb{R})\to C_{0}(G;\mathbb{R})$ is a linear map given by $\partial([u,v])=v-u$. The 1-homology group is defined as the kernel of the boundary map $H_{1}(G;\mathbb{R})=\mathrm{ker}(\partial)$. Its dimension $g=\dim(H_{1}(G;\mathbb{R}))$ is called the genus of $G$ and satisfies $g=m-n+1$. Any element $\sigma\in H_{1}(G;\mathbb{R})$ is called a 1-cycle. Let $\sigma(e)$ be the coefficient of $e$ in $\sigma$. The support of a 1-cycle is the set $\{e\in E:\sigma(e)\neq 0\}$, which forms a cycle in the sense of subgraph.
Next, we describe how to represent the graph as a matrix in terms of 1-cycles. Let $\ell:E\to\mathbb{R}_{+}$ be a length function. Define an inner product $\mathcal{Q}_{G}$ on $C_{1}(G;\mathbb{R})$ by assigning $\mathcal{Q}_{G}(e_{i},e_{j})=\delta_{ij}\sqrt{\ell(e_{i})\ell(e_{j})}$ on the edges and extending bilinearly to the whole space $C_{1}(G;\mathbb{R})$. Notice that under the inner product $\mathcal{Q}_{G}$, an element $e_{i}\in E$ has norm $\sqrt{\ell(e_{i})}$ when viewed as a 1-chain in $C_{1}(G;\mathbb{R})$, in contrast to its length $\ell(e_{i})$ when viewed as an edge in $G$ [14]. The inner product $\mathcal{Q}_{G}$ is compatible with edge subdivision: if $e$ is subdivided into $e^{\prime}$ and $e^{\prime\prime}$, then
@@
The 1-homology group $H_{1}(G;\mathbb{R})$ inherits the inner product $\mathcal{Q}_{G}$ as a closed subspace of $C_{1}(G;\mathbb{R})$. Fix a 1-cycle basis $\sigma_{1},\ldots,\sigma_{g}$ for $H_{1}(G;\mathbb{R})$. The inner product $\mathcal{Q}_{G}$ is represented by a matrix $Q$.
Definition 2.
Let $\mathcal{M}_{g}$ be the space of weighted graphs of genus $g$, and $\mathrm{SPD}(g)$ be the space of $g\times g$ SPD matrices. The tropical Torelli map is given by
$\mathcal{T}:\mathcal{M}_{g}\to\mathrm{SPD}(g),\,G\mapsto Q(G).$
@@ 3.2 Main Algorithm
For a connected graph $G$, we first compute a minimal spanning tree $T$. Fix an arbitrary orientation for $G$. For any edge $e=[u,v]\in G\backslash T$, it generates a 1-cycle $\sigma$ such that $\sigma=e+\sum_{e^{\prime}\in\gamma}\sigma(e^{\prime})e^{\prime},$ where $\gamma$ is the unique path in $T$ connecting $v$ to $u$. Since there are $m-(n-1)=g$ edges not in $T$, they generate a 1-cycle basis $\sigma_{1},\ldots,\sigma_{g}$ for $H_{1}(G;\mathbb{R})$. Construct a $g\times m$ matrix $M$ by setting $M_{ij}=\sigma_{i}(e_{j})$. The matrix $M$ is called the cycle–edge incidence matrix [7].
We reorder the edge set as follows: the first $n-1$ edges $e_{1},\ldots,e_{n-1}$ correspond to the minimal spanning tree $T$, while the remaining $g$ edges are sorted in ascending order by length. Then each 1-cycle $\sigma_{i}$ corresponds to the edge $e_{n-1+i}$. After reordering, the cycle–edge incidence matrix is in the block matrix form $M=[M_{T},I_{g}]$ where $I_{g}$ is the $g\times g$ identity matrix. Let $L=\mathrm{diag}\{\ell(e_{1}),\ldots,\ell(e_{m})\}$ be the diagonal matrix of edge lengths and $L_{T}=\mathrm{diag}\{\ell(e_{1}),\ldots,\ell(e_{n-1})\}$ and $L_{g}=\{\ell(e_{n}),\ldots,\ell(e_{m})\}$ be its blocks. Then the inner product $\mathcal{Q}_{G}$ on $H_{1}(G;\mathbb{R})$ is represented by the matrix
@@
1
2
Input: weighted graph $G$
Output: matrix $Q$
3 $Q\leftarrow[\,]$
4 for connected component $G_{i}=(E_{i},V_{i})$ do
5      compute a minimal spanning tree $T$
6      sort $E_{i}$ such that $e_{1},\ldots,e_{n_{i}-1}$ are in $T$, and $\ell(e_{n_{i}})<\ldots<\ell(e_{m_{i}})$
7      $M_{T}\leftarrow\mathrm{zeros}(g_{i},n_{i}-1)$
8      for $j\leftarrow 1$ to $g_{i}$  do
9           $[u,v]\leftarrow e_{n_{i}-1+j}$, find path $\gamma$ from $v$ to $u$ in $T$
10           for $k\leftarrow 1$ to $(n_{i}-1)$ do
11                $M_{jk}\leftarrow\gamma(e_{k})$
12
13                end for
14
15                end for
16               $Q\leftarrow\mathrm{diag}\{Q,M_{T}L_{T}M_{T}^{\top}+L_{g_{i}}\}$
17
18                end for
19

### v2

@@ Abstract
We introduce the first graph kernels for metric graphs via tropical algebraic geometry. In contrast to conventional graph kernels based on graph combinatorics such as nodes, edges, and subgraphs, our metric graph kernels are purely based on the geometry and topology of the underlying metric space. A key characterizing property of our construction is its invariance under edge subdivision, making the kernels intrinsically well-suited for comparing graphs representing different underlying metric spaces. We develop efficient algorithms to compute our kernels and analyze their complexity, which depends primarily on the genus of the input graphs rather than their size. Through experiments on synthetic data and selected real-world datasets, we demonstrate that our kernels capture complementary geometric and topological information overseen by standard combinatorial approaches, particularly in label-free settings. We further showcase their practical utility with an urban road network classification task.
Graph Kernels, Metric Graphs, Geometric Learning, Kernel Methods
@@ 1 Introduction
Graph kernels are powerful tools for measuring similarities between graphs and networks and have become essential in many machine learning tasks. Traditional graph kernels typically rely on combinatorial aspects of graphs, such as nodes, edges, and subgraphs, with additional information from node and edge labels and attributes (Kriege et al., 2020; Vishwanathan et al., 2010; Borgwardt et al., 2020; Nikolentzos et al., 2021; Neumann et al., 2016). While such kernels have demonstrated success in many domains, they can be limited in scenarios where graphs arise from geometric data, and where labels or attributes are absent or incomparable. Such challenges are intrinsic to metric graphs and render existing graph kernels inapplicable in this setting. In this paper, we introduce the first kernel framework designed specifically for metric graphs.
@@
Metric graphs are geometric realizations of graphs with a length function on their edges. To develop our metric graph kernels, we leverage the known one-to-one correspondence between metric graphs and abstract tropical curves from tropical algebraic geometry (Chan, 2021; Cao and Monod, 2025). This correspondence has been studied in tropical geometry via the tropical Torelli map (Chan, 2012; Brannetti et al., 2011), which sends any metric graph to a flat torus. In computational settings and especially machine learning, for a weighted graph with generic length function, we can compute a unique symmetric positive definite (SPD) matrix to represent the flat torus. Using the tropical Torelli map, we can effectively identify weighted graphs with SPD matrices, which contain the intrinsic geometric and topological information on the underlying metric graph.
@@
The space of SPD matrices is central in information geometry and can be endowed with various Riemannian metrics to capture its rich geometric structure (Luo et al., 2021; Zhang et al., 2019; Thanwerdas and Pennec, 2023b). However, it is only applicable when the metric graphs are of the same genus. In practice, to compare metric graphs of different genus via the tropical Torelli map, we must embed the SPD matrices into a common ambient space, namely the space of positive semi-definite (PSD) matrices. Motivated by the geometric structure of the tropical Jacobian, we seek a distance on the PSD space that is invariant under orthogonal transformations, which turns out to coincide with the well-known Bures–Wasserstein distance from information geometry (Bhatia et al., 2019; Thanwerdas and Pennec, 2023a). Leveraging this distance, we propose the tropical Torelli–Wasserstein (TTW) kernel for metric graphs. For simplicity and efficient computation, we also propose the tropical Torelli–Euclidean (TTE) kernel based on the Euclidean distance. Our graph kernels are purely based on the geometry and topology of the underlying metric graphs, thus are well-suited for comparing graphs that represent different underlying spaces.
@@
Contributions. Our work is the first to propose kernels for metric graphs, which we achieve by merging tropical geometry and information geometry in the context of machine learning. Our main contributions are the following:
@@
  • We propose new graph kernels for metric graphs, leveraging the tropical Torelli map from tropical geometry. Our graph kernels are invariant under edge subdivisions and naturally extend to well-defined kernels on the underlying space of metric graphs.
@@
  • We develop algorithms to compute our metric graph kernels. We show that the computational complexity is dominated by graph genus, a topological invariant that measures the number of independent cycles, rather than the number of nodes or edges. We give concrete analyses on different levels of graph sparsity.
  • We empirically evaluate our proposed kernels on synthetic and real-world datasets to assess their behavior in label-free settings. Our experiments illustrate that our kernels capture geometric and topological information that is complementary to existing combinatorial graph kernels, and are particularly effective when node and edge labels are unavailable.
  • We further demonstrate the practical relevance of our approach through targeted case studies. On synthetic datasets, we validate the theoretical runtime analysis, while on real-world data we apply our kernels to an urban road network classification task.
@@ 2 From Graphs to Metric Graphs
A graph consists of discrete sets of nodes and edges. A metric graph is a 1-dimensional metric space that can be realized as the underlying space of a graph. Any graph-based quantity extends well to metric graphs if it is compatible with edge subdivisions—a concept we will now introduce mathematically, which will then be motivated and illustrated concretely by road networks.
Edge Subdivisions. Let $G=(V,E)$ be a graph and $\ell:E\to\mathbb{R}_{+}$ be a length function. An edge subdivision of $e=[u,v]\in E$ is the following operation on $G$: First, add a new node $w$ to $V$, and then replace $e$ by two edges $e^{\prime}=[u,w]$ and $e^{\prime\prime}=[w,v]$ such that $\ell(e)=\ell(e^{\prime})+\ell(e^{\prime\prime})$. An edge subdivision increases both the number of nodes and number of edges by 1. $G^{\prime}$ is called a refinement of $G$, denoted as $G^{\prime}\geq G$, if $G^{\prime}$ can be obtained from $G$ by a sequence of edge subdivisions. By definition, all refinements of $G$ share the same underlying metric graph $|G|$.
Note that refinement introduces a new relation between graphs that differs from the usual notion of subgraph inclusion: If $G^{\prime}$ is a refinement of $G$, then there exists an injection $V(G)\to V(G^{\prime})$ and a surjection $E(G^{\prime})\to E(G)$. However, if $G$ is a subgraph of $G^{\prime}$, then $V(G)\to V(G^{\prime})$ and $E(G)\to E(G^{\prime})$ are both inclusions.
Metric Graph Kernels. Let $G_{1},G_{2}$ be two graphs. A graph kernel is a function whose value $k(x_{1},x_{2})$ quantifies the similarity between $G_{1},G_{2}$, typically in terms of their combinatorial structure. For example, the Weisfeiler–Lehman kernel assigns a large value when $G_{1}$ and $G_{2}$ are highly similar from the perspective of graph isomorphism (Shervashidze et al., 2009). In contrast, the following definition of a metric graph kernel aims to quantify similarity based on the underlying metric spaces. Specifically, the kernel $k(G_{1},G_{2})$ takes a large value when the metric realizations $|G_{1}|$ and $|G_{2}|$ are similar, for instance, from the perspective of isometry between metric spaces.
Definition 2.1.
A graph kernel $k$ is called a metric graph kernel if $k(G_{1}^{\prime},G_{2}^{\prime})=k(G_{1},G_{2})$ for any refinements $G_{1}^{\prime}\geq G_{1}$ and $G_{2}^{\prime}\geq G_{2}$, that is, the kernel $k$ is invariant under graph refinements.
Existing graph kernels are often based on nodes, edges, subgraphs, and node/edge labels, which are not invariant under graph refinements. Therefore, they fail to be metric graph kernels. In particular, node/edge labels are not well-defined on metric graphs—instead, the appropriate analogue should be functions on metric spaces. Table 1 summarizes comparisons of key notions between graphs and metric graphs.
Table 1: Key notions for graphs and their corresponding notions in the setting of metric graphs.
Graph Metric Graph
Vertex/Node Point
Edge Line segment
Subgraph Subspace
Label/Attribute Function
Graph distance Geodesic distance
Graph isomorphism Metric isometry
@@
3 The Tropical Torelli Map
In tropical geometry, metric graphs are known to be in one-to-one correspondence with abstract tropical curves (Mikhalkin and Zharkov, 2008). The tropical Torelli map is a fundamental construction that sends each tropical curve to a canonically defined flat torus, known as the tropical Jacobian variety. It serves as a tropical analogue of the classical Torelli map for algebraic curves, which sends a curve to its Jacobian variety. Further background and details on algebraic and tropical geometry are given in Appendix A.
The tropical Torelli map is originally defined in an intrinsic, basis-free manner via moduli spaces of tropical curves and principally polarized Abelian varieties (Brannetti et al., 2011). While this formulation is conceptually natural, it is not well suited for explicit computations. For computational purposes, it is necessary to instead express the map in a vector form by fixing bases for the relevant vector spaces and lattices of the tropical Jacobian varieties. This consideration motivates the main objective of this section. We first introduce the tropical Torelli map for weighted graphs; this is the map we use to send a graph to a matrix. For weighted graphs with generic length functions, we provide an algorithm for its practical computation, where the uniqueness of the resulting matrix representation is required but challenging to achieve. Under the generic assumption on length functions, our algorithm outputs a unique SPD matrix for any weighted graph.
@@ 3.1 Graph Homology
A concept that is central to our construction is the 1-homology group of a graph. The concept arises from algebraic topology and intuitively corresponds to cycles in a graph. We now give mathematical details on the construction of the 1-homology group of a graph and how we obtain a matrix representation of the graph in terms of 1-cycles via the tropical Torelli map.
We first outline the 1-homology group as follows. Let $G=(V,E)$ be a connected graph with $n$ nodes and $m$ edges. Let $C_{0}(G;\mathbb{R})$ be the $n$-dimensional vector space spanned by $V$ and $C_{1}(G;\mathbb{R})$ be the $m$-dimensional vector space spanned by $E$. Fix an arbitrary orientation on $G$. The boundary map $\partial:C_{1}(G;\mathbb{R})\to C_{0}(G;\mathbb{R})$ is a linear map given by $\partial([u,v])=v-u$. The 1-homology group is defined as the kernel of the boundary map $H_{1}(G;\mathbb{R})=\mathrm{ker}(\partial)$. Its dimension $g=\dim(H_{1}(G;\mathbb{R}))$ is called the genus of $G$ and satisfies g=mn+1g=m-n+1. Any element $\sigma\in H_{1}(G;\mathbb{R})$ is called a 1-cycle. Let $\sigma(e)$ be the coefficient of $e$ in $\sigma$. The support of a 1-cycle is the set $\{e\in E:\sigma(e)\neq 0\}$, which forms a cycle in the sense of subgraph.
Next, we describe how to represent the graph as a matrix in terms of 1-cycles. Let $\ell:E\to\mathbb{R}_{+}$ be a length function. Define an inner product $\mathcal{Q}_{G}$ on $C_{1}(G;\mathbb{R})$ by assigning $\mathcal{Q}_{G}(e_{i},e_{j})=\delta_{ij}\sqrt{\ell(e_{i})\ell(e_{j})}$ on the edges and extending bilinearly to the whole space $C_{1}(G;\mathbb{R})$. Notice that under the inner product $\mathcal{Q}_{G}$, an element $e_{i}\in E$ has norm $\sqrt{\ell(e_{i})}$ when viewed as a 1-chain in $C_{1}(G;\mathbb{R})$, in contrast to its length $\ell(e_{i})$ when viewed as an edge in $G$ (Ji, 2012). The inner product $\mathcal{Q}_{G}$ is compatible with edge subdivision: if $e$ is subdivided into $e^{\prime}$ and $e^{\prime\prime}$, then
@@
The 1-homology group $H_{1}(G;\mathbb{R})$ inherits the inner product $\mathcal{Q}_{G}$ as a closed subspace of $C_{1}(G;\mathbb{R})$. For any graph, choosing a 1-cycle basis $\sigma_{1},\ldots,\sigma_{g}$ for $H_{1}(G;\mathbb{R})$ is equivalent to choosing a spanning forest (Cao and Monod, 2025). Once a 1-cycle basis is fixed, the inner product $\mathcal{Q}_{G}$ is represented by a SPD matrix $Q$. The tropical Torelli map for weighted graphs sends graph $G$ to the matrix $Q(G)$.
@@ 3.2 Main Algorithm
For a connected graph $G$, we first compute a minimal spanning tree $T$. Fix an arbitrary orientation for $G$. For any edge e=[$\Te=[u,v]\in G\backslash$ T, it generates a 1-cycle $\sigma$ such that $\sigma=e+\sum_{e^{\prime}\in\gamma}\sigma(e^{\prime})e^{\prime},$ where $\gamma$ is the unique path in $T$ connecting $v$ to $u$. Since there are m(n1)=gm-(n-1)=g edges not in $T$, they generate a 1-cycle basis $\sigma_{1},\ldots,\sigma_{g}$ for $H_{1}(G;\mathbb{R})$. Construct a $g\times m$ matrix $M$ by setting $M_{ij}=\sigma_{i}(e_{j})$. The matrix $M$ is called the cycle–edge incidence matrix (Cao and Monod, 2025).
We reorder the edge set as follows: the first n1n-1 edges $e_{1},\ldots,e_{n-1}$ correspond to the minimal spanning tree $T$, while the remaining $g$ edges are sorted in ascending order by length. Then each 1-cycle $\sigma_{i}$ corresponds to the edge $e_{n-1+i}$. After reordering, the cycle–edge incidence matrix is in the block matrix form $M=[M_{T},I_{g}]$ where $I_{g}$ is the $g\times g$ identity matrix. Let $L=\mathrm{diag}\{\ell(e_{1}),\ldots,\ell(e_{m})\}$ be the diagonal matrix of edge lengths and $L_{T}=\mathrm{diag}\{\ell(e_{1}),\ldots,\ell(e_{n-1})\}$ and $L_{g}=\{\ell(e_{n}),\ldots,\ell(e_{m})\}$ be its blocks. Then the inner product $\mathcal{Q}_{G}$ on $H_{1}(G;\mathbb{R})$ is represented by the matrix
@@
0: weighted graph $G$
0: matrix $Q$
[…]
2:for each connected component $G_{i}=(E_{i},V_{i})$ do
3:  compute a minimal spanning tree $T$
4:  sort $E_{i}$ so that $e_{1},\ldots,e_{n_{i}-1}\in T$ and $\ell(e_{n_{i}})<\cdots<\ell(e_{m_{i}})$
5:  $M_{T}\leftarrow\mathrm{zeros}(g_{i},n_{i}-1)$
6:  for $j=1$ to $g_{i}$ do
7:   $u,v]\leftarrow e_{n_{i}-1+j}$
8:   find path $\gamma$ from $v$ to $u$ in $T$
9:   for $k=1$ to $n_{i}-1$ do
10:    $M_{jk}\leftarrow\gamma(e_{k})$
11:   end for
12:  end for
13:  $Q\leftarrow\mathrm{diag}\{Q,\,M_{T}L_{T}M_{T}^{\top}+L_{g_{i}}\}$
14:end for